# Every Pixel Tells a Story: End-to-End Urdu Newspaper OCR

### v1

@@ Abstract
This paper introduces a comprehensive end-to-end pipeline for Optical Character Recognition (OCR) on Urdu newspapers. In our approach, we address the unique challenges of complex multi-column layouts, low-resolution archival scans, and diverse font styles. Our process decomposes the OCR task into four key modules: (1) article segmentation, (2) image super-resolution, (3) column segmentation, and (4) text recognition. For article segmentation, we fine-tune and evaluate YOLOv11x to identify and separate individual articles from cluttered layouts. Our model achieves a precision of 0.963 and mAP@50 of 0.975. For super-resolution, we fine-tune and benchmark the SwinIR model (reaching 32.71 dB PSNR) to enhance the quality of degraded newspaper scans. To do our column segmentation, we use YOLOv11x to separate columns in text to further enhance performance - this model reaches a precision of 0.970 and mAP@50 of 0.975. In the text recognition stage, we benchmark a range of LLMs from different families, including Gemini, GPT, Llama, and Claude. The lowest WER of 0.133 is achieved by Gemini-2.5-Pro.
Every Pixel Tells a Story: End-to-End Urdu Newspaper OCR
@@ 1 Introduction
OCR plays a critical role in the digitization of printed material, enabling large-scale archival, search, and accessibility applications across languages and domains (Adenekan, 2024). In particular, newspaper OCR is essential for preserving records, supporting digital journalism, and making large volumes of textual data machine-readable. However, OCR for newspapers presents unique challenges due to complex layouts, variable font styles, and poor print quality.
These challenges are further compounded in low-resource languages like Urdu. Urdu script is inherently cursive, with context-sensitive letter forms, frequent ligatures, and baseline variations that make character segmentation and recognition difficult (Zaki et al., 2019). Urdu newspapers introduce an additional layer of complexity with multi-article, multi-column cluttered layouts, and stylized fonts, all of which degrade the performance of conventional OCR pipelines.
This paper introduces an end-to-end pipeline consisting of (1) Image Segmentation (article and column based), (2) Image Super Resolution, and (3) Text Recognition for OCR of Urdu newspapers, addressing the aforementioned challenges. Our contributions can be summarized as follows:
  1. 1.
    We fine-tune two separate YOLOv11x (Redmon et al., 2016) models: one for article-level segmentation and another for column-level segmentation, addressing the multi-column structure of Urdu newspaper layouts.
  2. 2.
    We fine-tune SwinIR (Liang et al., 2021) for image enhancement, improving degraded archival scans via super-resolution.
  3. 3.
    We benchmark LLMs including Gemini-2.5-Pro (Google et al., 2024), GPT-4o and 4.1 (OpenAI et al., 2024), Llama-4-Scout and Maverick (Meta, ), and Claude-3.7-Sonnet (Anthropic, ) for Urdu text recognition.
  4. 4.
    We release the dataset for LLM benchmarking and the first publicly available models for article and column segmentation for newspapers, and super-resolution for Urdu text  [^1] https://github.com/paper-seven/UrduOCR.
@@ 2 Related Work
Figure 1: End-to-end OCR pipeline with article and column segmentation, super-resolution, and LLM-based text recognition.
@@ 2.1 Prior Work in Urdu OCR
Early work in Urdu OCR focused on creating segmentation-based approaches such as those by Sabbour and Shafait (2013) who reported an accuracy of 91% on clean Urdu text, and Tabassam et al. (2009) who developed an OCR system for Urdu Naskh font utilizing pattern matching techniques, achieving an accuracy of 89%. However, these methods often struggled with the cursive and context-sensitive nature of the Urdu script.
To address these challenges, segmentation-free approaches were explored. Shabbir and Siddiqi (2016) proposed a system that recognizes Urdu words in the Nastaliq font using ligatures as units of recognition, employing Hidden Markov Models (HMMs) for classification. Similarly, Khattak (2017a) presented a segmentation-free OCR system for printed Urdu text, also using HMMs and focusing on ligature recognition. These approaches demonstrated improved handling of the complex ligature structures inherent in the Urdu script.
@@ 3.1 Datasets
Table 1 summarizes the dataset sizes used for each stage of our pipeline. The OCR evaluation dataset comprises 829 samples and includes 8,785 unique words and 2,539 sentences, demonstrating significant lexical diversity. All transcriptions were generated manually and verified by a second annotator to ensure high accuracy. Unlike traditional OCR datasets that focus on sentence-level inputs, our samples consist of full paragraphs of Urdu text to enable rigorous evaluation of LLMs for text recognition. For the image super-resolution dataset, we generated low-quality inputs by downscaling high-resolution Urdu newspaper scans by a factor of 4 and reducing the quality by 30% using JPEG compression.
Stage Train Size Test Size
@@
Text Recognition - 829
@@ 3.2 Pipeline
In the first stage, we use YOLOv11x for article-level segmentation. Following this, we apply our super-resolution model to the cropped article images in stage two. By enhancing image clarity and sharpening text details, this step produces cleaner, more uniform text blocks, which in turn enables more accurate column segmentation and more accurate text recognition. The low-quality cropped articles are passed through SwinIR model fine-tuned to increase the image resolution by 4 times. In the third stage we introduce column segmentation following the observation of improved model performance on shorter, single-column text blocks compared to full multi-column articles. This additional step breaks down complex layouts into more manageable units, improving recognition accuracy. In the fourth and final stage, we used LLMs to extract the text from each high-quality image segment.
@@ 3.3 Experimental Setup
In this paper, we fine-tune and evaluate YOLOv11x model for article and column level segmentation using standard object detection metrics including precision, recall, mAP@50, and mAP@50:95. The SwinIR model is trained and then evaluated using Peak Signal-to-Noise Ratio (PSNR) to measure reconstruction quality. To quantify the impact of super-resolution, we run our LLMs on both low- and high-resolution images generated by our model, comparing the word error rate (WER) and the character error rate (CER) between the two. For the final OCR stage, we assess LLM performance using WER and CER against human-annotated transcriptions. The system prompt provided to the LLMs is included below.
You are an OCR system. Your job is to transcribe image text exactly as shown, without interpretation, paraphrasing, translation, summarization, or hallucination.
@@ 4.1 Image Segmentation
For article-level segmentation, our fine-tuned YOLOv11x delivers both high accuracy and coverage. It achieves a precision of 0.963—meaning over 96% of its detected regions are true articles—and a recall of 0.971, capturing more than 97% of all actual article blocks. It also scores an mAP@50 of 0.975 and an mAP@50:95 of 0.866, demonstrating robust detection under both lenient and strict IoU thresholds. Together, these metrics confirm that our model reliably isolates article regions from the complex, multi-column layouts of Urdu newspapers, providing a solid foundation for super-resolution and column segmentation.
@@
Our fine-tuned YOLOv11x demonstrates outstanding performance on column-level segmentation. The precision of 0.970 indicates that 97.0% of all predicted column regions correspond to actual columns, minimizing false positives. The recall of 0.997 indicates that 99.7% of true column regions are successfully detected, ensuring almost no columns are missed. Under a loose IoU threshold (mAP@50), the model achieves 0.975, reflecting high tolerance for localization variation. Even when evaluated under the stricter mAP@50:95 metric, it maintains a robust score of 0.854, showing consistent accuracy across tighter overlap requirements. These results confirm that the model can reliably identify and localize column segments in complex Urdu newspaper layouts, providing a solid foundation for subsequent text extraction tasks.
@@ 4.2 Image Super-Resolution
After 100 epochs of training article-segmented data, SwinIR achieves a 32.71 dB PSNR on the test set. This level of image reconstruction is noteworthy given the challenges posed by Urdu script. A PSNR above 30 dB is generally considered strong for image restoration tasks (Kim et al., 2025; Anandhi and Mahalingam, 2025). In the context of Urdu OCR, it shows that the textual feature such as stroke continuity and glyph boundaries are effectively preserved which is essential for accurate downstream text recognition. Appendix A presents visual examples of low- and high-resolution Urdu newspaper images, demonstrating the improvements achieved by our model.
As shown in Table 3, our super-resolution model significantly improves text recognition performance across LLMs. Notably, Gemini-2.5-Pro, which already performs well on low-resolution inputs with a WER of 0.177, still benefits from enhancement, reducing its WER further to 0.133. Both GPT-4.1 and GPT-4o also show marked improvements, with WERs dropping from 0.682 to 0.254 and 0.779 to 0.327, respectively. Llama-4-Maverick also exhibit dramatic gains, decreasing WER from 1.036 to 0.305 after super-resolution. Claude-3.7-Sonnet and Llama-4-Scout fail to process most of the low-resolution images, often returning messages such as "Unfortunately, I am unable to extract text from the image…", "I can’t directly extract text…when the image quality does not allow for clear text recognition.", and "The image contains text in what appears to be Urdu script…" instead of performing any transcription.
4.3 Text Recognition
We evaluate text recognition performance of LLMs on the super-resolved images using WER and CER, as reported in Table 3.
Low-Res High-Res
LLM WER CER WER CER
GPT-4.1 0.682 0.471 0.254 0.096
GPT-4o 0.779 0.559 0.327 0.154
Gemini-2.5-Pro 0.177 0.046 0.133 0.032
Claude-3.7-Sonnet Fail Fail 0.249 0.100
Llama-4-Maverick 1.036 0.837 0.305 0.128
Llama-4-Scout Fail Fail 0.347 0.164
Each model was prompted using the following user message to ensure accurate text recognition without any unwanted additions:
Extract the exact text from this image. Preserve sentence structure NOT spacing. If anything is unreadable, write ’[UNREADABLE]’.
For high-resolution images, Gemini-2.5-Pro outperforms all other models, achieving the lowest WER and CER of 0.133 and 0.032, respectively. This indicates strong performance not only at the word level but also in character-level precision, making it well-suited for accurately transcribing Urdu. GPT-4.1 ranks as a close second, with a WER of 0.254 and CER of 0.096. While Claude-3.7-Sonnet struggles in low-resolution settings, it performs well on high-resolution images, achieving a WER of 0.249 and CER of 0.100, suggesting that its performance benefits significantly from our super-resolution module.
In contrast, GPT-4o underperforms with a WER of 0.327 and CER of 0.154, possibly due to overfitting to English structured text formats and reduced robustness for Urdu. Both Llama-4 models show relatively weaker performance, with Maverick outperforming Scout. Maverick records a WER of 0.305 and CER of 0.128, while Scout lags behind with a WER of 0.347 and CER of 0.164, aligning with expectations given Maverick’s larger model size.
5 Conclusion
Our proposed pipeline enables a mess-to-text transformation: complex, low-quality Urdu newspaper images are converted into clean, human-readable text without any manual correction or user intervention. By integrating fine-tuned models for article and column segmentation with super-resolution enhancements, we significantly improve the quality of inputs and therefore the results of large language models for OCR tasks.
These improvements lead to substantial performance gains in text recognition, significantly reducing WER and CER across models. Notably, they enable models like Claude-3.7-Sonnet, which previously failed under low-resolution conditions, to complete the OCR task successfully. This highlights the importance of layout-aware preprocessing and image enhancement for effective LLM-based approaches in low-resource settings.
Overall, we present a robust, modular system for high-quality Urdu OCR that significantly enhances the performance of general-purpose LLMs on complex scripts. Our pipeline demonstrates how targeted preprocessing can unlock capabilities of general purpose models for low-resource languages like Urdu in such tasks. We also release a curated text recognition evaluation dataset to support further research in this area.
6 Future Work
Moving forward, we plan to (i) expand our Urdu text recognition corpus and fine-tune open-source LLMs such as Gemma-3-12B and Gemma-3-27B on this data; (ii) develop OCR post-processing pipelines using LLMs to automatically correct residual errors and normalize transcriptions; and (iii) extend our evaluation of LLM-based text recognition to additional regional scripts—specifically Punjabi, Sindhi, and Balochi—to assess cross-language generalization and identify language-specific adaptation needs.
Limitations
Our evaluation dataset is relatively small, only 829 manually transcribed Urdu newspaper samples, which limits the statistical power and diversity of our OCR benchmarks (human annotation makes scaling difficult). We rely on off-the-shelf LLMs in a zero-shot setting; fine-tuning these models on in-domain OCR data could yield substantial gains but requires additional annotated text. Finally, the end-to-end pipeline incurs significant computational overhead (multiple YOLOv11x models, SwinIR training, and LLM inference), which may challenge deployment in resource-constrained environments.
Ethical Considerations
Appendix A Super-Resolution Results
Figure 2 shows qualitative results from our super-resolution model. Each row displays a pair of input and output images of our super-resolution model.

### v2

# From Press to Pixels: Evolving Urdu Text Recognition

@@ Abstract
This paper introduces an end-to-end pipeline for Optical Character Recognition (OCR) on Urdu newspapers, addressing challenges posed by complex multi-column layouts, low-resolution scans, and the stylistic variability of the Nastaliq script. Our system comprises four modules: (1) article segmentation, (2) image super-resolution, (3) column segmentation, and (4) text recognition. We fine-tune YOLOv11x for segmentation, achieving 0.963 precision for articles and 0.970 for columns. A SwinIR-based super-resolution model boosts LLM text recognition accuracy by 25–70%. We also introduce the Urdu Newspaper Benchmark (UNB), a manually annotated dataset for Urdu OCR. Using UNB and the OpenITI corpus, we compare traditional CNN+RNN-based OCR models with modern LLMs. Gemini-2.5-Pro achieves the best performance with a WER of 0.133. We further analyze LLM outputs via insertion, deletion, and substitution error breakdowns, as well as character-level confusion analysis. Finally, we show that fine-tuning on just 500 samples yields a 6.13% WER improvement, highlighting the adaptability of LLMs for Urdu OCR.
From Press to Pixels: Evolving Urdu Text Recognition
@@ 1 Introduction
OCR plays a critical role in the digitization of printed material, enabling large-scale archival, search, and accessibility applications across languages and domains (Adenekan, 2024). In particular, newspaper OCR is essential for preserving records, supporting digital journalism, and making large volumes of textual data machine-readable. OCR also facilitates broader accessibility when paired with systems such as Text-to-Speech models, it can make printed content available to visually impaired users or those with reading difficulties. However, OCR for newspapers presents unique challenges due to complex layouts, variable font styles, and poor print quality.
These challenges are further compounded in low-resource languages like Urdu, where the script is inherently cursive, with context-sensitive letter forms, frequent ligatures, and baseline variations that make segmentation and recognition difficult (Zaki et al., 2019). Urdu newspapers introduce an additional layer of complexity with multi-article, multi-column cluttered layouts, and stylized fonts, all of which further degrade the performance of conventional OCR pipelines. A particularly important factor in Urdu OCR is the typographic distinction between the Naskh and Nastaliq scripts. While most OCR systems, including Tesseract (Smith, 2007), EasyOCR (AI, 2020), Kraken (Kiessling, 2020), and UTRNet (Rahman et al., 2023) achieve reasonable performance on Naskh, they struggle significantly with Nastaliq script. The models trained on Urdu book dataset may perform adequately on data from the same domain but fail to generalize to other real-world settings like newspapers.
To address this, we introduce an end-to-end OCR pipeline for Urdu newspapers. Our approach incorporates layout-aware preprocessing, image super-resolution, and LLM-based text recognition. As part of this work, we introduce a new manually annotated dataset of Urdu newspaper scans in Nastaliq script, which we release publicly to support future research. Alongside the dataset, we also release the first super-resolution model tailored for Urdu text, as well as our fine-tuned models for article and column segmentation [^1] https://github.com/paper-seven/UrduOCR. To support our pipeline, we conduct a comprehensive evaluation of both traditional OCR systems and LLM-based methods across Naskh and Nastaliq scripts, with a focus on their ability to generalize across domains such as books and newspapers. We also perform detailed error analysis including insertion, deletion, and substitution errors as well as character-level breakdowns to highlight where even LLM fail and to inform future modeling efforts. Finally, we fine-tune GPT-4o on a small subset of data to demonstrate the potential of lightweight, targeted adaptation for Urdu OCR.
@@ 2 Related Work
Figure 1: OCR pipeline with article and column segmentation, super-resolution, and LLM-based text recognition.
@@
2.1 Open-source OCR Engines
Tesseract is one of the most widely used OCR engines. It follows a traditional CNN+LSTM architecture and supports Urdu in its default language pack. EasyOCR uses a CRNN-based pipeline with CTC loss, providing a general-purpose OCR framework for over 80 languages, including Urdu. Kraken is a flexible OCR engine based on OCRopus and designed to support right-to-left and complex scripts. It allows custom model training and has been successfully applied to Arabic-script languages.
@@ 2.2 Prior Work in Urdu OCR
Early work in Urdu OCR focused on creating segmentation-based approaches such as those by Sabbour and Shafait (2013) who reported an accuracy of 91% on clean Urdu text, and Tabassam et al. (2009) who developed an OCR system for Urdu Naskh script utilizing pattern matching techniques, achieving an accuracy of 89%. However, these methods often struggled with the cursive and context-sensitive nature of the Urdu script.
To address these challenges, segmentation-free approaches were explored. Shabbir and Siddiqi (2016) proposed a system that recognizes Urdu words in the Nastaliq script using ligatures as units of recognition, employing Hidden Markov Models (HMMs) for classification. Similarly, Khattak (2017a) presented a segmentation-free OCR system for printed Urdu text, also using HMMs and focusing on ligature recognition. These approaches demonstrated improved handling of the complex ligature structures inherent in the Urdu script. UTRNet is a recent CNN+RNN-based model specifically designed for Urdu text recognition. It uses a deep convolutional encoder combined with recurrent and attention-based decoders to model complex ligatures in Nastaliq. Trained on a synthetically generated dataset and the UPTI corpus, UTRNet achieves state-of-the-art accuracy (92.97%) on its test set.
@@ 3.1 Datasets
[…]
To assess generalization across writing styles and domains, we also evaluate the models on the paragraph-level images from OpenITI corpus, [^2] https://github.com/OpenITI/arabic_print_data which contains printed Urdu text in both Naskh and Nastaliq scripts from books. This dual-source evaluation enables us to benchmark both traditional OCR models and LLMs in terms of their robustness to script variation and domain shift from books to newspaper content.
For article and column segmentation, we constructed training datasets using publicly available Urdu newspaper scans sourced from online archives. For the image super-resolution dataset, we generated low-quality inputs by downscaling high-resolution Urdu newspaper scans by a factor of 4 and reducing the quality by 30% using JPEG compression.
Dataset Train Size Test Size
@@
UNB (Nastaliq) - 829
OpenITI (Nastaliq) - 250
OpenITI (Naskh) - 250
@@ 3.2 Pipeline
In the first stage of our pipeline, we use a fine-tuned YOLOv11x (Redmon et al., 2016) model for article-level segmentation, which isolates individual articles from cluttered, multi-article newspaper layouts. This step is essential because performing column segmentation directly would make it impossible to preserve article boundaries because of the cluttered newspaper layout. In the second stage, we apply a super-resolution model, SwinIR (Liang et al., 2021) to the cropped article images. We perform this step before column segmentation because enhancing the visual quality produces sharper edges and better-defined text regions, which in turn improves the accuracy column segmentation. In the third stage, we perform column segmentation within each article using a separate YOLOv11x model. This further simplifies the layout by breaking complex, multi-column articles into manageable single-column text blocks, which improves OCR accuracy across models. Finally, in the fourth stage, we pass the high-quality column segments to an LLM for text recognition.
@@ 3.3 Experimental Setup
Our experimental pipeline is divided into four main components: segmentation, super-resolution, text recognition, and error analysis. Below, we detail the setup for each stage and the metrics used.
3.3.1 Segmentation
We fine-tune and evaluate the YOLOv11x model for both article- and column-level segmentation using standard object detection metrics: precision, recall, mAP@50, and mAP@50:95.
3.3.2 Super-Resolution
The SwinIR model is trained and evaluated using Peak Signal-to-Noise Ratio (PSNR) to measure image reconstruction quality. To assess the impact of super-resolution on OCR performance, we run each LLM on both the original low-resolution images and their enhanced high-resolution counterparts generated by SwinIR. We then compare the Word Error Rate (WER) and Character Error Rate (CER) across these two conditions to quantify the benefit of super-resolution. Additionally, we run our column segmentation model on both versions of the images and measure performance using mAP@50. This allows us to quantify how super-resolution improves the reliability of layout detection when applied prior to column segmentation.
3.3.3 Text Recognition
We report text recognition results using standard OCR metrics: WER and CER. Our evaluation is divided into two parts: (i) baseline OCR models including Tesseract, EasyOCR, Kraken, and UTRNet, and (ii) LLMs including Gemini-2.5-Pro (Google et al., 2024), GPT-4.1 and GPT-4o (OpenAI et al., 2024), Claude-3.7-Sonne (Anthropic, ), and Llama-4-Maverick and Llama-4-Scout (Meta, )
Baselines Models.
We conduct a two-part evaluation of traditional OCR systems to assess their performance across different scripts and domains. First, we evaluate these models on the OpenITI book dataset, separately for Naskh and Nastaliq scripts, to highlight that traditional CNN- and RNN-based OCR models struggle to perform well on Urdu despite being fine-tuned specifically for both Naskh and Nastaliq script. Next, we test the same models on our UNB newspaper dataset to demonstrate that models fine-tuned on Nastaliq data from one source often fail to generalize to data from other domains.
Modern LLMs.
We assess the performance of multiple LLMs on both the OpenITI dataset and our UNB dataset. This evaluation allows us to measure the effectiveness of LLMs for Urdu OCR and their ability to generalize across different scripts and content types. To further demonstrate the potential of LLMs in this domain, we fine-tune GPT on a small subset of 500 annotated images from the UNB dataset and evaluate the model on the remaining 329 test images. By comparing the model’s performance before and after fine-tuning, we show that even limited in-domain supervision can lead to substantial improvements highlighting the adaptability of LLMs to the complexities of Nastaliq script. Each model is evaluated with a temperature of 0 to ensure reproducibility, using the following system prompt and user message to minimize inconsistencies and guide model behavior.
System Prompt  You are an OCR system. Your job is to transcribe image text exactly as shown, without interpretation, paraphrasing, translation, summarization, or hallucination. User Message  Extract the exact text from this image. Preserve sentence structure NOT spacing. If anything is unreadable, write ’[UNREADABLE]’.
GPT Finetuning.
To explore the adaptability of large language models for Urdu OCR, we fine-tune GPT-4o on a small subset of 500 high-resolution article images from our UNB dataset for 2 epochs. The goal of this experiment is to assess whether minimal supervision can significantly improve recognition accuracy in a low-resource setting. We use the same prompting setup as in our zero-shot evaluations and fine-tune the model to better align its visual recognition capabilities with the structural and stylistic features of Nastaliq script found in newspapers. The model is then evaluated on the remaining 329 test samples.
@@ 4.1 Image Segmentation
For article-level segmentation, our fine-tuned YOLOv11x delivers both high accuracy and coverage. It achieves a precision of 0.963—meaning over 96% of its detected regions are true articles—and a recall of 0.971, capturing more than 97% of all actual article blocks. It also scores an mAP@50 of 0.975 and an mAP@50:95 of 0.866, demonstrating robust detection under both lenient and strict IoU thresholds. Together, these metrics confirm that our model reliably isolates article regions from the complex, multi-column layouts of Urdu newspapers.
@@
Our fine-tuned YOLOv11x demonstrates outstanding performance on column-level segmentation. The precision of 0.970 indicates that 97.0% of all predicted column regions correspond to actual columns, minimizing false positives. The recall of 0.997 indicates that 99.7% of true column regions are successfully detected, ensuring almost no columns are missed. Under a loose IoU threshold (mAP@50), the model achieves 0.975, reflecting high tolerance for localization variation. Even when evaluated under the stricter mAP@50:95 metric, it maintains a robust score of 0.854, showing consistent accuracy across tighter overlap requirements. These results confirm that the model can reliably identify and localize column segments in complex Urdu newspaper layouts.
We perform column segmentation before text recognition to address challenges posed by the multi-column layout of Urdu newspapers. Without this step, LLMs tend to extract text across columns non-sequentially, resulting in jumbled and incoherent output. This misordering disrupts sentence structure and readability. Figure 2 illustrates a reading path followed by an LLM when extracting text directly from a multi-column article image, highlighting the need for explicit column segmentation.
Figure 2: The sequence of text extracted by the LLMs from multi-column image.
@@ 4.2 Image Super-Resolution
After 100 epochs of training article-segmented data, SwinIR achieves a 32.71 dB PSNR on the test set. This level of image reconstruction is noteworthy given the challenges posed by Urdu script. A PSNR above 30 dB is generally considered strong for image restoration tasks (Kim et al., 2025; Anandhi and Mahalingam, 2025). In the context of Urdu OCR, it shows that the textual feature such as stroke continuity and glyph boundaries are effectively preserved which is essential for accurate downstream text recognition. Figure 3 presents visual examples of low- and high-resolution Urdu newspaper images, demonstrating the improvements achieved by our model.